OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN!!!!!!!!!!!!!!! Peter drove 1000 mi in 4 hours. At that rate, how far would he drive in 3.5 hours? @Littlebird

OpenStudy (danjs):

Distance = rate x time 1000 mi = rate x 4 hrs Find Rate- then Distance = (that rate) x 3.5 hrs

OpenStudy (danjs):

solve 1000 = 4*r r = 1000/4 = 250 His rate r is 250 miles per hour So in 3.5 hours , he drives..... Distance = Rate x Time = 250 x 3.5 = ........

OpenStudy (danjs):

this exact question is in the closed.. hmm what

OpenStudy (anonymous):

250x3.5=875 @DanJS

OpenStudy (danjs):

right

OpenStudy (danjs):

All you have to know for this problem is Distance = Rate x Time

OpenStudy (danjs):

Given Distance = 1000 Time = 4 hours Distance = Rate x Time 1000 = Rate x 4 Solve for rate, and it is 250 Same Formula again Distance = Rate x Time Distance = 250 x 3.5 = 875

OpenStudy (anonymous):

so its 875 miles right????

OpenStudy (anonymous):

Lol yes.

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

I fanned u at least